OpenStudy (anonymous):

Find all real solutions of the equation. (Enter your answers as a comma-separated list. If there is no real solution, enter NO REAL SOLUTION.) x + 42=x

OpenStudy (anonymous):

dont forget to close your question once you are satisfied with an answer

OpenStudy (anonymous):

NOT THE ANSWER

OpenStudy (anonymous):

\[x+42=x\]\[0+42=0 ~~~\text{X}\] \[x+42=x\]\[~~~~~~42=0~~~\text{X}\] the answer \(is\) No Solution